OpenStudy (anonymous):

Solve the equation for m: A = 2(m - w). A. m = (A + w)/2 B. m = A + w/2 C. m = 2(A + w) D. m = A/2 + w

OpenStudy (anonymous):

Are you sure those are the answer choices? i keep getting m=a+2w/2

OpenStudy (anonymous):

yes its online

jigglypuff314 (jigglypuff314):

to solve for m in the equation of A = 2(m - w) first distribute so A= 2m - 2w then minus both sides by 2w so 2m = A + 2w then divide both sides by 2 so m = (A+2w)/2